The human body has an average specific heat capacity of 3500 J /(kg °C). If a 65 kg physics student drinks a can of coke (200 Cal ; 1Cal = 4186 J ) and all of its energy is converted to thermal energy , by how much would body temperature increase?


1
Expert's answer
2021-01-24T19:06:05-0500

The amount of energy required to rise the temperature of the body of mass "m = 65kg" by "\Delta T" is given with the following formula:


"Q =cm\Delta T"

where "c = 3500J/(kg\cdot \degree C)" is the specific heat capacity. Since "Q = 200cal = 200\cdot 4186J = 837200J", obtain:


"\Delta T = \dfrac{Q}{cm}\\
\Delta T = \dfrac{837200}{3500\cdot 65} = 3.68\degree C"

Answer. "3.68\degree C".